A student uses a launcher to fire Nerf darts horizontally from a height of 1.0 m and measures the distance the darts travel before they land. The average horizontal range of the darts is 6.5 m. Estimate the time the darts were in the air and from this, and estimate the launch velocity


1
Expert's answer
2021-07-22T10:02:44-0400

"h=\frac{gt^2}2,\implies"

"t=\sqrt{\frac{2h}g}=0.45~s,"

"v=\frac lt=14~\frac ms."